# An adaptation of InfoMap to absorbing random walks using absorption-scaled graphs

### v3

@@ 4.1.2 Main results for the map function $L^{(a)}$
Suppose that $P_{\delta}=P_{l}(D_{\delta},I,1)=(I-W(W+D_{\delta})^{-1})+A(W+D_{\delta})^{-1}=D_{\vec{r}}+Q$. Let $\pi_{\delta}^{(na)}$ be the stationary distribution of $P_{\delta}$, and let $\pi_{0}\mathrel{\mathop{:}}=D_{\vec{t}}D_{\vec{r}}\pi_{\delta}^{(na)}=(\pi_{\delta,i}^{(na)}t_{i}\delta_{i}/(\omega_{i}+\delta_{i}))_{i}$. We have that
@@ 4.2 Relating $\tilde{G}(D_{\delta},I)$ to $\tilde{G}(D_{\delta},\mathbf{0})$
Let $P_{0}=AW^{-1}$ be the transition-probability matrix of the discrete-time Markov chain that is associated with ~$\tilde{\mathcal{L}}(D_{\delta},\mathbf{0})$, and let $P_{1}=(A+D_{\delta})(W+D_{\delta})^{-1}$ be the transition-probability matrix of the discrete-time Markov chain that is associated with ~$\tilde{\mathcal{L}}(D_{\delta},I)$. Let $\pi$ and $\pi^{\prime}$ be the stationary distributions that are associated with $P_{0}$ and $P_{1}$, respectively. Let $Z_{i}$ be the fundamental matrix that is associated with $P_{i}$ (with $i\in\{1,2\}$). Let $U\mathrel{\mathop{:}}=(1/(\vec{\delta}^{\operatorname{\,T}}\vec{u}))\vec{u}\vec{1}^{\operatorname{\,T}}$ and $\alpha\mathrel{\mathop{:}}=\vec{\delta}^{\operatorname{\,T}}\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u}+\vec{\delta}^{\operatorname{\,T}}\vec{u})$, where $\vec{u}=(u_{1},\ldots,u_{n})^{\operatorname{\,T}}$ is a vector in the kernel $\operatorname{Ker}{(W-A)}$ with positive entries $u_{i}$ such that $\sum_{i=1}^{n}u_{i}=1$. We have that
@@
(Absorption inverse) Let $-~\tilde{\mathcal{L}}$ be the unnormalized graph Laplacian matrix of a strongly connected graph, and let $\vec{d}$ be an absorption-rate vector. Let $D\mathrel{\mathop{:}}=\operatorname{diag}\{\vec{d}\}$, $N_{1,0}\mathrel{\mathop{:}}=\{\vec{x}\in\mathbb{R}^{n}\mathrel{\mathop{:}}D\vec{x}\in\operatorname{Range}\tilde{\mathcal{L}}\}$, and $R_{1,0}\mathrel{\mathop{:}}=\{D\vec{x}\mathrel{\mathop{:}}\vec{x}\in\operatorname{Ker}\tilde{\mathcal{L}}\}$. An absorption inverse ~$\tilde{\mathcal{L}}^{\vec{d}}$ of $-~\tilde{\mathcal{L}}$ with respect to $\vec{d}$ is defined by the following properties:
@@
Let $-~\tilde{\mathcal{L}}$ be the unnormalized graph Laplacian matrix of a strongly connected graph, let $\vec{d}$ be an absorption-rate vector, and let $D=\operatorname{diag}\{\vec{d}\}$. The following relationship holds:
@@
Let ~$\tilde{\mathcal{L}}^{\vec{d}}$ be the absorption inverse of $-~\tilde{\mathcal{L}}$ with respect to $\vec{d}$. Let $\vec{u}$ be a vector in Ker~normal-$-~\operatorname{Ker}\tilde{\mathcal{L}}$ with positive entries $u_{i}$ such that $\sum_{i}u_{i}=1$. Additionally, let $D\mathrel{\mathop{:}}=\operatorname{diag}\{\vec{d}\}$ and $U\mathrel{\mathop{:}}=(1/\hat{\delta})\vec{u}\vec{1}^{\operatorname{\,T}}$, where $\hat{d}=\sum_{i}(u_{i}d_{i})$. The following relationship holds:
@@
For $\epsilon\mathrel{\mathop{:}}={\lVert\mathcal{L}^{\vec{\delta}}D_{\delta}\rVert}\ll 1$, it follows from (33) that
@@
Let $\vec{d}$ be an absorption-rate vector, and let $\vec{u}\in\operatorname{Ker}\tilde{\mathcal{L}}=\operatorname{Ker}{(W-A)}$. Additionally, let $\vec{w}\mathrel{\mathop{:}}=W_{\operatorname{diag}}$, $\pi\mathrel{\mathop{:}}=W\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u})$, $D\mathrel{\mathop{:}}=\operatorname{diag}\{\vec{\delta}\}$, $U\mathrel{\mathop{:}}=(1/(\vec{d}^{\operatorname{\,T}}\vec{u}))\vec{u}\vec{1}^{\operatorname{\,T}}$, and $Z\mathrel{\mathop{:}}=W^{-1}Z_{0}$, where $Z_{0}=\left(I-AW^{-1}+\pi\vec{1}^{\operatorname{\,T}}\right)^{-1}$ is the fundamental matrix that is associated with $AW^{-1}$. We have that
@@
Let ~$\tilde{\mathcal{L}}_{1}\mathrel{\mathop{:}}=\tilde{\mathcal{L}}(D_{\delta},I)=(W-A)(W+D_{\delta})^{-1}$, and let $\vec{d}_{1}$ be the diagonal of $D_{\delta}(W+D_{\delta})^{-1}$. Additionally, let $U\mathrel{\mathop{:}}=\vec{u}\vec{1}^{\operatorname{\,T}}/(\vec{\delta}^{\operatorname{\,T}}\vec{u})$, $U_{1}\mathrel{\mathop{:}}=(W+D_{\delta})U$, and $D_{1}\mathrel{\mathop{:}}=D_{\delta}(W+D_{\delta})^{-1}$. We have that
@@
Let $\vec{d}^{\prime}\mathrel{\mathop{:}}=\vec{d}_{\mathrm{s}}(D_{\delta},I)$ be the absorption-rate vector that is associated with $H=I$. The following proposition relates the absorption inverse $\mathcal{L}^{\vec{\delta}}$ (which is associated with $\tilde{G}(D_{\delta},\mathbf{0})$) and the absorption inverse $\mathcal{L}^{\vec{d}^{\prime}}$ (which is associated with $\tilde{G}(D_{\delta},I)$). We prove this proposition in Appendix A.
@@
Let $\vec{d}^{\prime}\mathrel{\mathop{:}}=\vec{d}_{\mathrm{s}}(D_{\delta},I)=\vec{w}+\vec{\delta}=(\omega_{1}+\delta_{1},\ldots,\omega_{n}+\delta_{n})^{\operatorname{\,T}}$ be the scaled rate vector that is associated with the absorption-scaled graph $\tilde{G}(D_{\delta},I)$. With $\mathcal{L}=W-A$, $\alpha\mathrel{\mathop{:}}=\vec{\delta}^{\operatorname{\,T}}\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u}+\vec{\delta}^{\operatorname{\,T}}\vec{u})$, $\pi=W\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u})$, $Z_{0}=\left(I-AW^{-1}+\pi\vec{1}^{\operatorname{\,T}}\right)^{-1}$, and $Z_{*}=W^{-1}\left(Z_{0}-\pi\vec{1}^{\operatorname{\,T}}\right)$, it follows that
@@ 5.1 A three-node network
Figure 4: An example three-node network. The relatively large size of node 2 indicates that its node-absorption rate is larger than the absorption rates of the other nodes.
@@ 5.2 A four-clique network
Consider a network that consists of four planted cliques [see Figure 6(a)], with four nodes each. Additionally, suppose that all of the edge weights are $1$. The node-absorption rates of the nodes in the cliques $C_{1}\mathrel{\mathop{:}}=\{1,2,3,4\}$ and $C_{3}\mathrel{\mathop{:}}=\{9,10,11,12\}$ are $\delta_{i}=7$ (with $i\in C_{1}\cup C_{3}$), and the node-absorption rates of the nodes in the cliques $C_{2}\mathrel{\mathop{:}}=\{5,6,7,8\}$ and $C_{4}\mathrel{\mathop{:}}=\{13,14,15,16\}$ are $\delta_{i}=1$ (with $i\in C_{2}\cup C_{4}$).
Consider the node-absorption-rate vector $\vec{\delta}=(\delta_{1},\ldots,\delta_{16})^{\operatorname{\,T}}$ and the absorption-scaled graphs $\tilde{G}(D_{\delta},\mathbf{0})$ and $\tilde{G}(D_{\delta},(3/2)I)$. We use Algorithm 1a with inputs $P_{l}(D_{\delta},\mathbf{0},t)$ and $P_{l}(D_{\delta},(3/2)I,t)$ and Algorithm 1b with inputs $P_{e}(D_{\delta},\mathbf{0},t)$ and $P_{e}(D_{\delta},(3/2)I,t)$ for different Markov times $t$. Arguably, the network structure (i.e., the network topology and the edge weights) on its own favors the partition $M^{*}\mathrel{\mathop{:}}=\{C_{1},C_{2},C_{3},C_{4}\}$ in Figure 6(b). However, the larger absorption rates in the cliques $C_{1}$ and $C_{3}$ lead to the partition $M^{**}\mathrel{\mathop{:}}=\{\{1\},\{2\},\{3\},\{4\},C_{2},\{9\},\{10\},\{11\},\{12\},C_{4}\}$ in Figure 6(c).
Figure 6: We show an example network with four planted cliques and two partitions of its set of nodes. In (a), the node-absorption rates of the large nodes are $\delta_{i}=7$ and the node-absorption rates of the small nodes are $\delta_{i}=1$. The edge weights are all equal to $1$. The node sizes are positively correlated with (but are not linearly related to) the node-absorption rates of the associated nodes. In (b) and (c), each color indicates a different community. The partition $M^{*}$ in (b) is the four-clique planted partition, which arises from the network structure (i.e., the network topology and the edge weights). The partition $M^{**}$ in (c) arises from a combination of the network structure and the node-absorption rates.
@@ 5.3 A square-lattice network with four different node-absorption rates
We consider a square-lattice network with 36 nodes (see Figure 8). We divide the set of nodes into four sublattices (with labels $B_{1}$, $B_{2}$, $B_{3}$, and $B_{4}$), which we illustrate using nodes of different sizes in Figure 8. We endow the nodes in sublattice $B_{1}$ with a node-absorption rate of $0.2$, the nodes in sublattice $B_{2}$ with a node-absorption rate of $0.7$, the nodes in sublattice $B_{3}$ with a node-absorption rate of $1.5$, and the nodes in sublattice $B_{4}$ with a node-absorption rate of $1.7$.
Figure 8: A square-lattice network with nodes that have node-absorption rates in the set $\{0.2,0.7,1.2,1.7\}$. The node sizes are positively correlated with (but are not linearly related to) the node-absorption rates of the associated nodes.
@@
Figure 9: (a,b) The numbers of communities in the partitions that we obtain using (a) Algorithm 1a with the input $P_{l}(D_{\delta},\mathbf{0},t)$ and (b) Algorithm 1b with the input $P_{e}(D_{\delta},I,t)$. In (c) and (d), we show the partitions that we obtain from Algorithm 1a with the input $P_{l}(D_{\delta},\mathbf{0},0.04)$ and Algorithm 1b with the input $P_{e}(D_{\delta},I,5.25)$, respectively. The colors in panels (c) and (d) indicate community assignments; the node sizes are positively correlated with (but are not linearly related to) the node-absorption rates of the associated nodes.
@@ 6.1 An example of Salathé and Jones [38]
Salathé and Jones [38] explored the effect of changes in community structure on the outbreak duration, final size, and outbreak peak in simulations of an SIR model on networks. In our work, we consider a modified version of one of the examples in [38]. We describe their example in several stages. Each stage $s$ has an associated network $G^{\mathrm{WS}}_{s}$ and SIR simulations on that network.
@@
In stage $s=1$, Salathé and Jones [38] considered a network $G^{\mathrm{WS}}_{1}$ that consists of Watts–Strogatz (WS) small-world graphs [32], which yield “planted communities” that are connected to each other by “community bridges” through edges that are assigned uniformly at random from all possible node pairs. The set of nodes is the same in all stages, but Salathé and Jones rewire the set of edges in each stage. In stage $s+1$, Salathé and Jones rewired one community bridge of $G^{\mathrm{WS}}_{s}$ (which is known from stage $s$) into one edge within a planted community in the following manner. First, they selected a community bridge $\{i_{1},i_{2}\}$ uniformly at random from all of the community bridges of $G^{\mathrm{WS}}_{s}$. They then selected a node $i_{k}$ uniformly at random from $\{i_{1},i_{2}\}$ and selected a node $i_{3}\neq i_{k}$ uniformly at random from the planted community of $i_{k}$. Finally, they obtained the set of edges of $G^{\mathrm{WS}}_{s+1}$ by removing $\{i_{1},i_{2}\}$ from the union of the edge $\{i_{k},i_{3}\}$ and the set of edges of $G^{\mathrm{WS}}_{s}$. At each stage $s$, they simulated an SIR process on $G^{\mathrm{WS}}_{s}$ and recorded the means of the outbreak durations, final sizes, and outbreak peaks. In all stages, they considered the same time-independent transmission and recovery intensities in their SIR simulations.
@@
We also describe the stochastic SIR model that Salathé and Jones employed in [38]. We simulate this SIR model on networks using a Gillespie algorithm  [2, 13, 22]. Each node is either in the susceptible state (i.e., compartment) $S$, the infected state $I$, or the recovered state $R$. If a susceptible node has a single infected neighbor, then the time that the node remains susceptible is exponentially distributed with transmission intensity [^6] Our use of the word “intensity” follows the terminology in [2]. $\beta$. With this distribution, the mean time is $1/\beta$. The time that an infected node remains infected is exponentially distributed with recovery intensity $\gamma$, which yields a mean time of $1/\gamma$. Let $|{\rm S}|$ denote the number of susceptible nodes, $|{\rm I}|$ denote the number of infected nodes, and $|{\rm R}|$ denote the number of recovered nodes. Additionally, $|{\rm SI}|$ denotes the number of neighboring node pairs in which one node is susceptible and the others node is infected. This SIR model is a continuous-time Markov chain with two possible events: (1) one susceptible node is infected at a rate $\beta|{\rm SI}|$ or (2) one infected node recovers at a rate $\gamma|{\rm I}|$. Two events cannot occur simultaneously. In Table 2, we show the transitions in the numbers of nodes in each state for this model.

### v4

@@ 4.1.2 Main results for the map function $L^{(a)}$
Suppose that $P_{\delta}=P_{l}(D_{\delta},I,1)=(I-W(W+D_{\delta})^{-1})+A(W+D_{\delta})^{-1}=D_{\vec{r}}+Q$. Let $\pi_{\delta}^{(na)}$ be the stationary distribution of $P_{\delta}$, and let $\pi_{0}:=D_{\vec{t}}D_{\vec{r}}\pi_{\delta}^{(na)}=(\pi_{\delta,i}^{(na)}t_{i}\delta_{i}/(\omega_{i}+\delta_{i}))_{i}$. We have that
@@ 4.2 Relating $\tilde{G}(D_{\delta},I)$ to $\tilde{G}(D_{\delta},\mathbf{0})$
Let $P_{0}=AW^{-1}$ be the transition-probability matrix of the discrete-time Markov chain that is associated with ~$\tilde{\mathcal{L}}(D_{\delta},\mathbf{0})$, and let $P_{1}=(A+D_{\delta})(W+D_{\delta})^{-1}$ be the transition-probability matrix of the discrete-time Markov chain that is associated with ~$\tilde{\mathcal{L}}(D_{\delta},I)$. Let $\pi$ and $\pi^{\prime}$ be the stationary distributions that are associated with $P_{0}$ and $P_{1}$, respectively. Let $Z_{i}$ be the fundamental matrix that is associated with $P_{i}$ (with $i\in\{1,2\}$). Let $U:=(1/(\vec{\delta}^{\operatorname{\,T}}\vec{u}))\vec{u}\vec{1}^{\operatorname{\,T}}$ and $\alpha:=\vec{\delta}^{\operatorname{\,T}}\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u}+\vec{\delta}^{\operatorname{\,T}}\vec{u})$, where $\vec{u}=(u_{1},\ldots,u_{n})^{\operatorname{\,T}}$ is a vector in the kernel $\operatorname{Ker}{(W-A)}$ with positive entries $u_{i}$ such that $\sum_{i=1}^{n}u_{i}=1$. We have that
@@
(Absorption inverse) Let $\tilde{\mathcal{L}}$ be the unnormalized graph Laplacian matrix of a strongly connected graph, and let $\vec{d}$ be an absorption-rate vector. Let $D:=\operatorname{diag}\{\vec{d}\}$, $N_{1,0}:=\{\vec{x}\in\mathbb{R}^{n}:D\vec{x}\in\operatorname{Range}\tilde{\mathcal{L}}\}$, and $R_{1,0}:=\{D\vec{x}:\vec{x}\in\operatorname{Ker}\tilde{\mathcal{L}}\}$. An absorption inverse ~$\tilde{\mathcal{L}}^{\vec{d}}$ of $\tilde{\mathcal{L}}$ with respect to $\vec{d}$ is defined by the following properties:
@@
Let $\tilde{\mathcal{L}}$ be the unnormalized graph Laplacian matrix of a strongly connected graph, let $\vec{d}$ be an absorption-rate vector, and let $D=\operatorname{diag}\{\vec{d}\}$. The following relationship holds:
@@
Let ~$\tilde{\mathcal{L}}^{\vec{d}}$ be the absorption inverse of $\tilde{\mathcal{L}}$ with respect to $\vec{d}$. Let $\vec{u}$ be a vector in $\operatorname{Ker}\tilde{\mathcal{L}}$ with positive entries $u_{i}$ such that $\sum_{i}u_{i}=1$. Additionally, let $D:=\operatorname{diag}\{\vec{d}\}$ and $U:=(1/\hat{\delta})\vec{u}\vec{1}^{\operatorname{\,T}}$, where $\hat{d}=\sum_{i}(u_{i}d_{i})$. The following relationship holds:
@@
For $\epsilon:={\lVert\mathcal{L}^{\vec{\delta}}D_{\delta}\rVert}\ll 1$, it follows from (33) that
@@
Let $\vec{d}$ be an absorption-rate vector, and let $\vec{u}\in\operatorname{Ker}\tilde{\mathcal{L}}=\operatorname{Ker}{(W-A)}$. Additionally, let $\vec{w}:=W_{\operatorname{diag}}$, $\pi:=W\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u})$, $D:=\operatorname{diag}\{\vec{\delta}\}$, $U:=(1/(\vec{d}^{\operatorname{\,T}}\vec{u}))\vec{u}\vec{1}^{\operatorname{\,T}}$, and $Z:=W^{-1}Z_{0}$, where $Z_{0}=\left(I-AW^{-1}+\pi\vec{1}^{\operatorname{\,T}}\right)^{-1}$ is the fundamental matrix that is associated with $AW^{-1}$. We have that
@@
Let ~$\tilde{\mathcal{L}}_{1}:=\tilde{\mathcal{L}}(D_{\delta},I)=(W-A)(W+D_{\delta})^{-1}$, and let $\vec{d}_{1}$ be the diagonal of $D_{\delta}(W+D_{\delta})^{-1}$. Additionally, let $U:=\vec{u}\vec{1}^{\operatorname{\,T}}/(\vec{\delta}^{\operatorname{\,T}}\vec{u})$, $U_{1}:=(W+D_{\delta})U$, and $D_{1}:=D_{\delta}(W+D_{\delta})^{-1}$. We have that
@@
Let $\vec{d}^{\prime}:=\vec{d}_{\mathrm{s}}(D_{\delta},I)$ be the absorption-rate vector that is associated with $H=I$. The following proposition relates the absorption inverse $\mathcal{L}^{\vec{\delta}}$ (which is associated with $\tilde{G}(D_{\delta},\mathbf{0})$) and the absorption inverse $\mathcal{L}^{\vec{d}^{\prime}}$ (which is associated with $\tilde{G}(D_{\delta},I)$). We prove this proposition in Appendix A.
@@
Let $\vec{d}^{\prime}:=\vec{d}_{\mathrm{s}}(D_{\delta},I)=\vec{w}+\vec{\delta}=(\omega_{1}+\delta_{1},\ldots,\omega_{n}+\delta_{n})^{\operatorname{\,T}}$ be the scaled rate vector that is associated with the absorption-scaled graph $\tilde{G}(D_{\delta},I)$. With $\mathcal{L}=W-A$, $\alpha:=\vec{\delta}^{\operatorname{\,T}}\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u}+\vec{\delta}^{\operatorname{\,T}}\vec{u})$, $\pi=W\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u})$, $Z_{0}=\left(I-AW^{-1}+\pi\vec{1}^{\operatorname{\,T}}\right)^{-1}$, and $Z_{*}=W^{-1}\left(Z_{0}-\pi\vec{1}^{\operatorname{\,T}}\right)$, it follows that
@@ 5.1 A three-node network
Figure 4: An example three-node network. The node-absorption rate of node 2 is greater than or equal to the node-absorption rates of nodes 1 and 3.
@@ 5.2 A four-clique network
Consider a network that consists of four planted cliques [see Figure 6(a)], with four nodes each. Additionally, suppose that all of the edge weights are $1$. The node-absorption rates of the nodes in the cliques $C_{1}:=\{1,2,3,4\}$ and $C_{3}:=\{9,10,11,12\}$ are $\delta_{i}=7$ (with $i\in C_{1}\cup C_{3}$), and the node-absorption rates of the nodes in the cliques $C_{2}:=\{5,6,7,8\}$ and $C_{4}:=\{13,14,15,16\}$ are $\delta_{i}=1$ (with $i\in C_{2}\cup C_{4}$).
Consider the node-absorption-rate vector $\vec{\delta}=(\delta_{1},\ldots,\delta_{16})^{\operatorname{\,T}}$ and the absorption-scaled graphs $\tilde{G}(D_{\delta},\mathbf{0})$ and $\tilde{G}(D_{\delta},(3/2)I)$. We use Algorithm 1a with inputs $P_{l}(D_{\delta},\mathbf{0},t)$ and $P_{l}(D_{\delta},(3/2)I,t)$ and Algorithm 1b with inputs $P_{e}(D_{\delta},\mathbf{0},t)$ and $P_{e}(D_{\delta},(3/2)I,t)$ for different Markov times $t$. Arguably, the network structure (i.e., the network topology and the edge weights) on its own favors the partition $M^{*}:=\{C_{1},C_{2},C_{3},C_{4}\}$ in Figure 6(b). However, the larger absorption rates in the cliques $C_{1}$ and $C_{3}$ lead to the partition $M^{**}:=\{\{1\},\{2\},\{3\},\{4\},C_{2},\{9\},\{10\},\{11\},\{12\},C_{4}\}$ in Figure 6(c).
Figure 6: An example network with four planted cliques and two partitions of its set of nodes. In (a), the node-absorption rates of the large nodes are $\delta_{i}=7$ and the node-absorption rates of the small nodes are $\delta_{i}=1$. The edge weights are all equal to $1$. In (b) and (c), each color indicates a different community. The partition $M^{*}$ in (b) is the four-clique planted partition, which arises from the network structure (i.e., the network topology and the edge weights). The partition $M^{**}$ in (c) arises from a combination of the network structure and the node-absorption rates.
@@ 5.3 A square-lattice network with four different node-absorption rates
We consider a square-lattice network with 36 nodes (see Figure 8). We divide the set of nodes into four sublattices (with labels $B_{1}$, $B_{2}$, $B_{3}$, and $B_{4}$), which we illustrate using nodes of different sizes in Figure 8. We endow the nodes in sublattice $B_{1}$ with a node-absorption rate of $0.2$, the nodes in sublattice $B_{2}$ with a node-absorption rate of $0.7$, the nodes in sublattice $B_{3}$ with a node-absorption rate of $1.2$, and the nodes in sublattice $B_{4}$ with a node-absorption rate of $1.7$.
Figure 8: A square-lattice network with nodes that have node-absorption rates in the set $\{0.2,0.7,1.2,1.7\}$.
@@
Figure 9: (a,b) The numbers of communities in the partitions that we obtain using (a) Algorithm 1a with the input $P_{l}(D_{\delta},\mathbf{0},t)$ and (b) Algorithm 1b with the input $P_{e}(D_{\delta},I,t)$. In (c) and (d), we show the partitions that we obtain from Algorithm 1a with the input $P_{l}(D_{\delta},\mathbf{0},0.04)$ and Algorithm 1b with the input $P_{e}(D_{\delta},I,5.25)$, respectively. The colors in panels (c) and (d) indicate community assignments.
@@ 6.1 An example of Salathé and Jones [38]
Salathé and Jones [38] explored the effect of differences in community structure on the outbreak duration, final size, and outbreak peak in simulations of an SIR model on networks. In our work, we consider a modified version of one of the examples in [38]. Salathé and Jones rewired edges to consider networks with different community structures, and they thereby examined how community structure impacts various aspects of disease outbreaks. Inspired by their exploration, we examine different node-absorption rates (instead of rewiring edges) to consider different community structures, and we simulate disease dynamics on the resulting networks. One of the findings of Salathé and Jones is that outbreak duration depends non-monotonically on how modular the network is. We examine whether there is a similar relationship between outbreak duration and the community structure that we detect with our adaptation of InfoMap.
@@
We begin by describing the example of Salathé and Jones. Each stage $s$ of their example has an associated network $G^{\mathrm{WS}}_{s}$ and SIR simulations on that network. In stage $s=1$, Salathé and Jones [38] considered a network $G^{\mathrm{WS}}_{1}$ that consists of Watts–Strogatz (WS) small-world graphs [32], which yield “planted communities” that are connected to each other by “community bridges” through edges that are assigned uniformly at random from all possible node pairs. The set of nodes is the same in all stages, but Salathé and Jones rewire the set of edges in each stage. In stage $s+1$, Salathé and Jones rewired one community bridge of $G^{\mathrm{WS}}_{s}$ (which is known from stage $s$) into one edge within a planted community in the following manner. First, they selected a community bridge $\{i_{1},i_{2}\}$ uniformly at random from all of the community bridges of $G^{\mathrm{WS}}_{s}$. They then selected a node $i_{k}$ uniformly at random from $\{i_{1},i_{2}\}$ and selected a node $i_{3}\neq i_{k}$ uniformly at random from the planted community of $i_{k}$. Finally, they obtained the set of edges of $G^{\mathrm{WS}}_{s+1}$ by removing $\{i_{1},i_{2}\}$ from the union of the edge $\{i_{k},i_{3}\}$ and the set of edges of $G^{\mathrm{WS}}_{s}$. At each stage $s$, they simulated an SIR process on $G^{\mathrm{WS}}_{s}$ and recorded the means of the outbreak durations, final sizes, and outbreak peaks. In all stages, they considered the same time-independent transmission and recovery intensities in their SIR simulations.
@@
We also describe the stochastic SIR model that Salathé and Jones employed in [38]. We simulate this SIR model on networks using a Gillespie algorithm  [2, 13, 22]. Each node is either in the susceptible state (i.e., compartment) $S$, the infected state $I$, or the recovered state $R$. If a susceptible node has $k$ infected neighbors, then the time that it remains susceptible is exponentially distributed with transmission intensity [^6] Our use of the word “intensity” follows the terminology in [2]. $\beta k$. With this distribution, the mean time is $1/(\beta k)$. The time that an infected node remains infected is exponentially distributed with recovery intensity $\gamma$, which yields a mean time of $1/\gamma$. Let $|{\rm S}|$ denote the number of susceptible nodes, $|{\rm I}|$ denote the number of infected nodes, and $|{\rm R}|$ denote the number of recovered nodes. Additionally, $|{\rm SI}|$ denotes the number of neighboring node pairs in which one node is susceptible and the others node is infected. This SIR model is a continuous-time Markov chain with two possible events: (1) one susceptible node is infected at a rate $\beta|{\rm SI}|$ or (2) one infected node recovers at a rate $\gamma|{\rm I}|$. Two events cannot occur simultaneously. In Table 2, we show the transitions in the numbers of nodes in each state for this model.